OpenStudy (anonymous):

Find the sum of the infinite geometric series: 4+1+1/4+1/16+1/64

OpenStudy (anonymous):

use \[S=\frac{a}{1-r}\] with \(a=4,r=\frac{1}{4}\)

OpenStudy (anonymous):

you can almost do it in your head \(1-\frac{1}{4}=\frac{3}{4}\) the reciprocal of \(\frac{3}{4}\) is \(\frac{4}{3}\) multiply by 4 you get \(\frac{16}{3}\)